OpenStudy (anonymous):

Find all x values at which the line tangent to the graph of the function f(x)=2sin^2(x) is horizontal

OpenStudy (anonymous):

Find the derivative and set it equal to zero! then find the x values!

OpenStudy (anonymous):

how do you do that? I'm soooo bad at this.

OpenStudy (anonymous):

\[2\sin ^{2}(x) \prime = 4\sin x \cos x\] Now set \[0 = 4 \sin x \cos x\] Which means either \[0 = 4\] or \[0=sin x\] or \[0 = \cos x \] Now get the values for x from those equations

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

find the values of x where the graphs of sin(x) and cosine(x) is zero! For example: sin(180) = 0 -> x = 180 is a solution! you will see a pattern